state what properties you use/defintions/etc...

Let (G,✳) be a group, and let a∈G. Let C(a) = {g∈G: a✳g = g✳a}.

In this problem we will prove that (C(a),✳) is a subgroup of (G,✳).

C(a)⊆G by the definition of C(a).
1
Expert's answer
2020-10-26T20:08:31-0400

Solution

"C(a)" is not empty since "\exist e \in G" such that "a *e=e*a" so "e\in C(a)" (where e is the identity element of G)

Since "C(a)\subseteq G" then "C(a)" is associative under the operation "*" .

Let "x,y\in C(a)", then "a*x=x*a" and "a*y=y*a"

"a*x*y=a*(x*y) =(x*y)*a" (by definition) so "(x*y)\in C(a)"

Since "e \in C(a)" we have "a*e=a*(x*x^{-1}) = (a*x)*x^{-1}" and "e*a=(x^{-1}*x)*a= x^{-1}*(x*a)" which implies that "x^{-1}\in C(a)"

Therefore "(C(a),*)" is a subgroup of the group "(G,*)" .